# Non-destructive imaging of bulk electrical ‘hidden’ state switching in a 1-TaS2 cryo-memory device

### v1

@@ V Switching-induced out-of-plane lattice contraction
As reported before for the o-HCDW state [46], we also observed a change in the out-of-plane lattice constant in the e-HCDW state. Figure S10 shows the 1D projection onto the out-of-plane reciprocal space coordinate l of the (004) lattice peak. The signal was measured on the entire flake (no spatial resolution) in the unswitched (A), the partially- (B) and the fully-switched state (C). The high intensity peak is the (004) lattice peak reflection. We note that because this measurement is not spatially-resolved, intensity (from the unswitched portions of the flake) still shows up there even in the C state. Clearly, in the B and C states a satellite lattice peak emerges from the HCDW state in the switching channel. Thus, we observed that for the e-HCDW state the out-of-plane lattice peak position shifts to a higher reciprocal space value, meaning that the lattice contracts by . The out-of-plane lattice constant difference between the switched and unswitched region leads to strain along the border which we speculate facilitates the switching in-plane but also in volume. Similarly, the c lattice constant contracts also between the NCCDW and CCDW states [69, 70, 71, 68].
We set two ROIs around the lattice signals in reciprocal space: and r.l.u. for the main and satellite peak, respectively, to obtain spatial maps by summing up the intensity within these ROIs (see Fig. S11). Also here we observe that the HCDW state switching channel appeared at the bottom left corner of the flake. Thus, not only the electronic CDW structure is modified in the e-HCDW state but also the out-of-plane lattice contracts in the non-volatile switching channel of the device. In the following, we show the agreement of the HCDW state switching channel observed from the out-of-plane lattice contraction and the HCDW peak.
As done for Fig. 3 of the main text, we define different spatial regions of the device: outside of the switching channel (light blue), near the switching channel (light purple) and in-between the electrodes (dark purple). We then map out the 2D reciprocal space projections of the (004) lattice peak (see Fig. S12). Unlike the CCDW peak which shows a characteristic elongation, the lattice peak is sharp also in the out-of-plane direction. The correlation of the ratio of the projected out-of-plane lattice intensities of the main and satellite signal, , and that of the HCDW and CCDW peak intensity, , is shown in Fig. S13a. Figure S13b compares the contours of the switching region based on the ratio of the lattice signals, and the HCDW/CCDW peaks, revealing that the two are also spatially correlated. Thus, also the lattice peak contraction can be used as a fingerprint to locate the HCDW switching.
@@ VI Finite element method simulations
where is the temperature, the volume and the heat source term. The density of the material , heat capacity [72], as well as the thermal [73] and electrical conductivity of all materials in the device were taken from literature or determined experimentally [20, 29]. The thermal model was coupled to the electromagnetic one which provided in the differential equations by calculating the resistive losses in the device, i.e. Joule heating.
@@
The structure and dimensions of the simulated device is shown in Fig. S14a. The geometry of the Au contacts and the 1-TaS2 crystal flake was realistically replicated both in shape and size, including irregularities such as the narrower spacing at one (left) side of the structure and the two additional disconnected contact fingers in the middle. Excitation of the device was applied via the two outer contacts by assuming square wave current pulses of amplitudes between and . At the beginning of each simulation, the device was fully in the unswitched state A at temperature. After applying the excitation, switching of the elements from A to the fully-switched state C was handled by a simplified model which predicts formation of the HCDW state domain wall network under the influence of a sufficiently large charge carrier injection. Thus, in the present simulations the state of each element was switched once the current density of the element reached a predefined current density threshold of , calibrated by the experiment. Figure S14b shows the switched material (orange) for different magnitudes of current excitation. All cases identify the preferential area of device switching at the left side of the device, which is influenced primarily by the narrower spacing between the contacts. With increasing current amplitude, the area gradually expands along the intergap space. Furthermore, we observe that the penetration depth of the switching is determined by the level of anisotropy of 1-TaS2. In the isotropic case (see top row of Fig. S14b), the switched volume propagates throughout the entire thickness of the crystal flake, whereas it remains closer to the surface when assuming an out/in-plane conductivity anisotropy of 1:5 (see middle row of Fig. S14b). A final set of simulations (see bottom row of Fig. S14b), we tested also the influence of the disconnected contact fingers. While they do affect the quantitative extent of the switched volume, the additional conductive paths they provide qualitative agree with the simulations in the top row of Fig. S14b.
@@
Contrary to the experiment these simulations do not reproduce the widening of the switched volume laterally and beneath the contacts. We speculate that this effect in the experiment is driven by propagation of local strain between the switched and unswitched area—a feature presently not included in the numerical simulations. We also observe that compared to the experiment, the device switches fully at a lower current amplitude (between to ). We attribute this to slight discrepancies between the assumed and effective material parameters. Finally, in Fig. S14c, device heating due to the current excitation is also demonstrated for the isotropic situation without influence of the disconnected middle contacts (top row in Fig. S14b). The uneven heating upon current pulse application localized to the left side of the device is also in agreement with the preferential path of the current flow across the 1-TaS2 flake.

### v2

@@ V Switching-induced out-of-plane lattice contraction
As reported before for the o-HCDW state [49], we also observed a change in the out-of-plane lattice constant in the e-HCDW state. FigureĀ 10 shows the 1D projection onto the out-of-plane reciprocal space coordinate l of the (004) lattice peak. The signal was measured on the entire flake (no spatial resolution) in the unswitchedĀ A, the partially- B and the fully-switched state C. The high intensity peak is the (004) lattice peak reflection. We note that because this measurement is not spatially-resolved, intensity (from the unswitched portions of the flake) still shows up there even in the C state. Clearly, in the B and C states a satellite lattice peak emerges from the HCDW state in the switching channel. Thus, we observed that for the e-HCDW state the out-of-plane lattice peak position shifts to a higher reciprocal space value, meaning that the lattice contracts by . The out-of-plane lattice constant difference between the switched and unswitched region leads to strain along the border which we speculate facilitates the switching in-plane but also in volume. Similarly, the c lattice constant contracts between the NCCDW and CCDW states [73, 74, 75, 72].
We set two ROIs around the lattice signals in reciprocal space: and r.l.u. for the main and satellite peak, respectively, to obtain spatial maps by summing up the intensity within these ROIs (Fig.Ā 11). Also here we observe that the HCDW state switching channel appeared at the bottom left corner of the flake. Thus, not only the electronic CDW structure is modified in the e-HCDW state but also the out-of-plane lattice contracts in the non-volatile switching channel of the device. In the following, we show the agreement of the HCDW state switching channel observed from the out-of-plane lattice contraction and the HCDW peak.
As done for Fig.Ā 3 of the main text, we define different spatial regions of the device: outside of the switching channel (light blue), near the switching channel (light purple) and in-between the electrodes (dark purple). We then map out the 2D reciprocal space projections of the (004) lattice peak (Fig.Ā 12). Unlike the CCDW peak which shows a characteristic elongation, the lattice peak is sharp also in the out-of-plane direction. The correlation of the ratio of the projected out-of-plane lattice intensities of the main and satellite signal, , and that of the HCDW and CCDW peak intensity, , is shown in Fig.Ā 13a. FigureĀ 13b compares the contours of the switching region based on the ratio of the lattice signals, and the HCDW/CCDW peaks, revealing that the two are also spatially correlated. Thus, also the lattice peak contraction can be used as a fingerprint to locate the HCDW switching.
@@ VI Finite element method simulations
where is the temperature, the volume and the heat source term. The density of the material , heat capacity [76], as well as the thermal [77] and electrical conductivity of all materials in the device were taken from literature or experiments [23, 32]. The thermal model was coupled to the electromagnetic one which provided in the differential equations by calculating the resistive losses in the device, i.e. Joule heating.
@@
The structure and dimensions of the simulated device is shown in Fig.Ā 14a. The geometry of the Au contacts and the 1-TaS2 crystal flake was realistically replicated both in shape and size, including irregularities such as the narrower spacing at one (left) side of the structure and the two additional disconnected contact fingers in the middle. Excitation of the device was applied via the two outer contacts by assuming square wave current pulses of amplitudes between and . At the beginning of each simulation, the device was fully in the unswitched state A at temperature. After applying the excitation, switching of the elements from A to the fully-switched state C was handled by a simplified model which predicts formation of the HCDW state domain wall network under the influence of a sufficiently large charge carrier injection. Thus, in the present simulations the state of each element was switched once the current density of the element reached a predefined current density threshold of , calibrated by the experiment. FigureĀ 14b shows the switched material (orange) for different magnitudes of current excitation. All cases identify the preferential area of device switching at the left side of the device, which is influenced primarily by the narrower spacing between the contacts. With increasing current amplitude, the area gradually expands along the intergap space. Furthermore, we observe that the penetration depth of the switching is determined by the level of anisotropy of 1-TaS2. In the isotropic case (Fig.Ā 14b, top row), the switched volume propagates throughout the entire thickness of the crystal flake, whereas it remains closer to the surface when assuming an out/in-plane conductivity anisotropy of 1:5 (Fig.Ā 14b, middle row). A final set of simulations (Fig.Ā 14b, bottom row), we tested also the influence of the disconnected contact fingers. While they do affect the quantitative extent of the switched volume, the additional conductive paths they provide qualitative agree with the simulations in the top row of Fig.Ā 14b.
@@
Contrary to the experiment these simulations do not reproduce the widening of the switched volume laterally and beneath the contacts. We speculate that this effect in the experiment is driven by propagation of local strain between the switched and unswitched areaāaĀ feature presently not included in the numerical simulations. We also observe that compared to the experiment, the device switches fully at a lower current amplitude (between to ). We attribute this to slight discrepancies between the assumed and effective material parameters. Finally, in Fig.Ā 14c, device heating due to the current excitation is also demonstrated for the isotropic situation without influence of the disconnected middle contacts (Fig.Ā 14b, top row). The uneven heating upon current pulse application localized to the left side of the device is also in agreement with the preferential path of the current flow across the 1-TaS2 flake.